50% H2O is in Na2SO4.XH2O then what is the value of X?
1
Expert's answer
2017-07-04T10:56:06-0400
Molar mass of Na2SO4 is
2∙23 + 32 + 4∙16 = 142
Molar mass of H2O is
2∙1 + 16 = 18
18X/142 =0.5/0.5=1
X =(142/18)≈8.
Answer: 8.